Good day

My question is:


What mass of KIO3 is needed to convert the copper in 0.3500g of CuSO4.5H2O to Cu(IO3)2


Please assist


1
Expert's answer
2021-05-08T23:53:59-0400

CuSO4.5H2O + 2KIO3 = Cu(IO3)2 + K2SO4 + 5H2O

n (KIO3) = 2 x n (CuSO4.5H2O)

n = m / M

M (CuSO4.5H2O) = 250 g/mol

n (CuSO4.5H2O) = 0.3500 / 250 = 0.0014 mol

n (KIO3) = 2 x 0.0014 = 0.0028 mol

M (KIO3) = 214 g/mol

m (KIO3) = n x M = 0.0028 x 214 = 0.6 g